# Visual Cues Enhance Predictive Turn-Taking for Two-Party Human Interaction

### v1

@@ Abstract
Turn-taking is richly multimodal. Predictive turn-taking models (PTTMs) facilitate naturalistic human-robot interaction, yet most rely solely on speech. We introduce MM-VAP, a multimodal PTTM which combines speech with visual cues including facial expression, head pose and gaze. We find that it outperforms the state-of-the-art audio-only in videoconferencing interactions (84% vs. 79% hold/shift prediction accuracy). Unlike prior work which aggregates all holds and shifts, we group by duration of silence between turns. This reveals that through the inclusion of visual features, MM-VAP outperforms a state-of-the-art audio-only turn-taking model across all durations of speaker transitions. We conduct a detailed ablation study, which reveals that facial expression features contribute the most to model performance. Thus, our working hypothesis is that when interlocutors can see one another, visual cues are vital for turn-taking and must therefore be included for accurate turn-taking prediction. We additionally validate the suitability of automatic speech alignment for PTTM training using telephone speech. This work represents the first comprehensive analysis of multimodal PTTMs. We discuss implications for future work and make all code publicly available.
@@ 1 Introduction
We introduce multimodal VAP (MM-VAP), a transformer-based PTTM which combines speech with visual features, including facial expression. We show our model outperforms the voice activity projection (VAP) model, a state-of-the-art audio-only PTTM Ekstedt and Skantze (2022b).
@@
In prior work, PTTM performance is reported as a single figure for all holds and shifts. In a more comprehensive analysis, we group holds and shifts by duration of silence between turns. This reveals the strength of our multimodal model across all durations of silence between turns (gaps) and overlapping speech. The performance of both MM-VAP and VAP is worse when there are longer gaps in the Candor corpus, however, MM-VAP outperforms the audio-only model across all durations of gaps considered (83% vs 79% for all gaps >0 ms, 78% vs 75% for all gaps >750 ms). We discuss our working hypothesis that interlocutors utilise visual cues when they can see one another, and they must therefore be included to maintain robust performance.
@@
We conclude with a discussion of our findings, highlighting the vital importance of non-verbal communication, which researchers in human-robot interaction must consider going forward. We make our code publicly available for future research [^1] https://github.com/russelsa/mm-vap.
@@ 2 Background
[…]
@@ 3 Methodology
We use OpenFace Baltrušaitis et al. (2016) to extract visual features. Facial action units (FAUs, 17 in total) numerically describe facial movements e.g. jaw drop on a scale from $0.0$ to $5.0$. There is one gaze vector per eye which is a $3$ dimensional unit vector. We also extract head position ($X,Y,Z$ in mm) and head rotation (roll, pitch and yaw in radians). We select 15 facial landmarks located on the brow, jaw, nose and lips ($x$ and $y$ in pixels) and a confidence score. In total, there is one $60$ dimensional vector per frame. We scale all features to 0, 1 by max min scaling at participant level. Histograms of eye gaze and head pose are unimodal with different modes per participant, reflecting the differing setup of participant devices. We therefore zero mean the head pose and eye gaze vectors at participant level. We show a sample visualisation of OpenFace features in Figure 1. Head pose is depicted as a blue cube centred on the head pointing in the estimated direction of head pose. Eye gaze vectors are depicted in green. Facial landmarks are shown in red, though we only use a subset of 15 in this work. It is not straightforward to visualise FAUs so these are omitted.
@@
We establish the performance of the state-of-the-art voice activity projection (VAP) model Ekstedt and Skantze (2022b), an audio-only turn-taking model which has been used in several studies e.g. Ekstedt and Skantze (2022a); Ekstedt et al. (2023); Inoue et al. (2024). The VAP model is a transformer Vaswani (2017) based neural network trained to predict who will speak in the next 2 seconds: the training objective, shown in Figure 3. At each point in time there are four bins per speaker representing the next 2 seconds. The bins span the next 0-200, 200-600, 600-1200 and 1200-2000 milliseconds. If 50% of frames within a bin contain speech, we set the bin label to 1, otherwise the bin is labelled 0. This gives $2562^{8}=$ 256 possible VAP states. The model has 5.8 M parameters, and full details are in Inoue et al. (2024); Ekstedt and Skantze (2022a).
@@
Like the VAP model, our model, multimodal VAP (MM-VAP, Figure 4) consists of self- and cross-attention blocks. A self-attention block is $N$ stacked transformer decoder layers, where query ($q$), key ($k$) and value ($v$) are identical:
[…]
A transformer layer learns attention or the ’compatibility’ of output with the query via the key and value Vaswani (2017). In self-attention blocks, as $q=k=v$, the model learns temporal patterns in the input e.g. audio or video. In cross-attention blocks, we stack $N$ layers with two inputs $x_{1}$ and $x_{2}$. In each layer, we compute two transformer layers with shared weights:
@@
[…]
@@
We conduct an ablation study by training 4 different versions of the MM-VAP model. We train and validate using the same procedure, but each of which receives audio and a subset of the visual features as input. We divide the visual features into gaze, head pose, facial action unit and facial landmarks groups (6, 6, 17, 30 dimensions receptively, excluding the scalar confidence score).
@@ 4.3 Multimodal turn-taking
We continue by investigating the performance of our proposed early and late fusion turn-taking models (MM-VAP). We train on the Candor corpus (Section 3). For completeness, we also train a VAP model only using video features. This model is identical to the original VAP model with the removal of the audio feature extractor and a dimensionality of 60, reflecting the dimensionality of the visual feature vector.
@@
(FTO > +250 ms) v 0.72 $\downarrow$140.82 $\downarrow$80.47 $\downarrow$3368 $\downarrow$15%
a+v (e) 0.84 - 0.90 - 0.74 $\uparrow$582 $\uparrow$3%
a+v (l) 0.86 $\uparrow$30.90 $\uparrow$20.74 $\uparrow$683 $\uparrow$4%
@@
(FTO > +250 ms) v 0.70 $\downarrow$130.80 $\downarrow$80.45 $\downarrow$3366 $\downarrow$14%
a+v (e) 0.83 $\uparrow$20.88 - 0.69 $\uparrow$479 $\uparrow$3%
a+v (l) 0.83 $\uparrow$30.89 $\uparrow$20.70 $\uparrow$680 $\uparrow$4%
@@
(FTO > 0 ms) v 0.78 $\downarrow$100.87 $\downarrow$50.43 $\downarrow$3570 $\downarrow$10%
a+v (e) 0.87 $\uparrow$20.92 - 0.69 $\uparrow$680 $\uparrow$3%
a+v (l) 0.87 $\uparrow$20.92 - 0.71 $\uparrow$683 $\uparrow$4%
@@
(FTO < -250 ms) v 0.72 $\downarrow$80.82 $\downarrow$30.42 $\downarrow$2564 $\downarrow$8%
a+v (e) 0.79 $\uparrow$30.87 $\uparrow$20.60 $\uparrow$772 $\uparrow$3%
a+v (l) 0.80 $\uparrow$40.87 $\uparrow$20.62 $\uparrow$1074 $\uparrow$5%
@@
We report the average performance over 5 folds of the Candor corpus in Table 3. As before, we consider model performance 200 ms before a shift/hold (FTO > 250 ms). Our multimodal models significantly outperform the audio-only and video-only models ($p<0.01$). The late fusion model shows a $3$% relative increase in the weighted $F_{1}$ score and a $6$% relative increase in balanced accuracy over the audio-only model. The video-only model has the worst overall performance with a 33% relative reduction in the $F_{1}$ shift score.
@@
Next, we remove the requirement for a minimum period of silence between turns, increasing the number of shifts and holds considered (FTO > 0 vs FTO > 250 ms in Table 1). We also move the evaluation window to 200 ms before the end of a turn, ensuring the next speaker has not started to speak during evaluation. This hence captures the capacity of the model to predict upcoming shifts while the previous speaker is still speaking, akin to human turn-taking (Section 1). For this more comprehensive set of speaker transitions, our multimodal models outperforms both video-only and audio-only models. Best performance is achieved with the late fusion model, with a $6\%$ relative increase in the $F_{1}$ shift score and a $4\%$ relative increase in balanced accuracy (Table 3). We also evaluate hold/shift prediction before overlapping speech. Note that for the purposes of this evaluation, we exclude overlapping speech where there is no change in speaker 1 second after the overlap has concluded. This ensures that brief periods of overlap are excluded (i.e. backchannels). We find a 6% increase in balanced accuracy and a $10\%$ increase in the $F_{1}$ shift scores for the late-fusion MM-VAP model (Table 3).
@@
The balanced accuracies in Tables 2 and 3 correspond with the 0 and 250 ms points in Figure 6. The gap in performance of Switchboard and Candor is consistent with our prior finding of a lower $F_{1}$ shift score on Switchboard (Table 2). Performance falls for both audio-only VAP and multimodal MM-VAP models as the duration of silence between speaking turns increases, indicating that these turns are more challenging to predict. However, we find that the performance of the MM-VAP models is significantly better than the audio-only VAP models across all FTOs ($p<0.01$ in all cases). Thus, the inclusion of visual information leads to improved performance across the complete range of holds and shifts in the corpus.
@@
Finally, we conduct an ablation study, training the late fusion version of the MM-VAP model on audio and gaze, head pose, facial landmarks and facial action unit subsets. We report the 5-fold average performance in Table 4.
@@
(FTO > +250 ms) a + gaze 0.74 $\downarrow$140.83 $\downarrow$80.51 $\downarrow$3167 $\downarrow$20%
a + pose 0.84 $\downarrow$20.89 - 0.72 $\downarrow$380 $\downarrow$3%
a + lmks 0.79 $\downarrow$80.86 $\downarrow$40.62 $\downarrow$1774 $\downarrow$11%
a + faus 0.85 - 0.89 - 0.73 $\downarrow$281 $\downarrow$2%
@@
(FTO > 250 ms) a + gaze 0.72 $\downarrow$130.82 $\downarrow$80.48 $\downarrow$3565 $\downarrow$19%
a + pose 0.82 - 0.87 - 0.68 $\downarrow$878 $\downarrow$2%
a + lmks 0.78 $\downarrow$70.85 $\downarrow$40.59 $\downarrow$2072 $\downarrow$10%
a + faus 0.83 - 0.88 - 0.69 $\downarrow$679 $\downarrow$2%
@@
(FTO > 0 ms) a + gaze 0.78 $\downarrow$100.88 $\downarrow$50.43 $\downarrow$3965 $\downarrow$20%
a + pose 0.87 - 0.92 - 0.67 $\downarrow$379 $\downarrow$2%
a + lmks 0.83 $\downarrow$50.90 $\downarrow$30.57 $\downarrow$1873 $\downarrow$10%
a + faus 0.87 - 0.92 - 0.68 $\downarrow$280 $\downarrow$4%
@@
(FTO < 250 ms) a + gaze 0.68 $\downarrow$150.81 $\downarrow$70.35 $\downarrow$4458 $\downarrow$22%
a + pose 0.79 $\downarrow$20.86 - 0.58 $\downarrow$671 $\downarrow$3%
a + lmks 0.74 $\downarrow$80.84 $\downarrow$40.49 $\downarrow$2266 $\downarrow$11%
a + faus 0.79 $\downarrow$20.86 - 0.59 $\downarrow$472 $\downarrow$2%
@@ 5 Discussion
Our re-implementation of the audio-only VAP PTTM Ekstedt and Skantze (2022b); Inoue et al. (2024) performed well on the Candor videoconferencing (VC) corpus. However, our new multimodal PTTM, MM-VAP - which uses speech along with facial action units, landmarks, gaze, and head pose - significantly outperformed the VAP model. The performance increase was most notable on the $F_{1}$ shift metric, with a 6-10% relative increase in performance (Table 2). Returning to our first research question, our findings show that visual cues improve PTTM performance. This echoes the psycholinguistics literature underlining the importance of visual cues in turn-taking (Section 1).
@@
MM-VAP has shown promising results, but it can be improved. OpenFace could be replaced by a more powerful learnable visual front-end, e.g. from the audio-visual speech recognition literature Fenghour et al. (2021). The model should be updated to handle audio and video at different sample rates (Section 3).

### v2

@@ Abstract
Turn-taking is richly multimodal. Predictive turn-taking models (PTTMs) facilitate naturalistic human-robot interaction, yet most rely solely on speech. We introduce $M$-VAP, a multimodal PTTM which combines speech with visual cues including facial expression, head pose and gaze. We find that it outperforms the state-of-the-art audio-only in videoconferencing interactions (84% vs. 79% hold/shift prediction accuracy). Unlike prior work which aggregates all holds and shifts, we group by duration of silence between turns. This reveals that through the inclusion of visual features, $M$-VAP outperforms a state-of-the-art audio-only turn-taking model across all durations of speaker transitions. We conduct a detailed ablation study, which reveals that facial expression features contribute the most to model performance. Thus, our working hypothesis is that when interlocutors can see one another, visual cues are vital for turn-taking and must therefore be included for accurate turn-taking prediction. We additionally validate the suitability of automatic speech alignment for PTTM training using telephone speech. This work represents the first comprehensive analysis of multimodal PTTMs. We discuss implications for future work and make all code publicly available.
@@ 1 Introduction
We introduce multimodal VAP ($M$-VAP), a transformer-based PTTM which combines speech with visual features, including facial expression. We show our model outperforms the voice activity projection (VAP) model, a state-of-the-art audio-only PTTM Ekstedt and Skantze (2022b).
@@
In prior work, PTTM performance is reported as a single figure for all holds and shifts. In a more comprehensive analysis, we group holds and shifts by duration of silence between turns. This reveals the strength of our multimodal model across all durations of silence between turns (gaps) and overlapping speech. The performance of both $M$-VAP and VAP is worse when there are longer gaps in the Candor corpus, however, $M$-VAP outperforms the audio-only model across all durations of gaps considered (83% vs 79% for all gaps >0 ms, 78% vs 75% for all gaps >750 ms). We discuss our working hypothesis that interlocutors utilise visual cues when they can see one another, and they must therefore be included to maintain robust performance.
@@
We conclude with a discussion of our findings, highlighting the vital importance of non-verbal communication, which researchers in human-robot interaction must consider going forward. We make our code publicly available for future research [^1] https://github.com/russelsa/$m$-vap.
@@ 2 Background
Turn-taking is aided by multimodal cues. The words that we speak as well as how we speak (prosody) play important roles. Bögels and Torreira (2015) found prosody is essential for listeners to correctly determine the end of the turns with multiple completion points e.g. "are you a student $\backslash$ here $\backslash$ at this university $\backslash$.". Visual cues include gaze. Speakers look away at the start and back toward the listener at the end of a turn Kendon (1967). Certain gestures and facial expressions are associated with faster shifts between speakers Trujillo et al. (2021); Nota et al. (2023).
@@ 3 Methodology
We use OpenFace Baltrušaitis et al. (2016) to extract visual features. Facial action units (FAUs, 17 in total) numerically describe facial movements e.g. jaw drop on a scale from $0.0$ to $5.0$. There is one gaze vector per eye which is a 33 dimensional unit vector. We also extract head position ($X,Y,Z$ in $m$) and head rotation (roll, pitch and yaw in radians). We select 15 facial landmarks located on the brow, jaw, nose and lips ($x$ and $y$ in pixels) and a confidence score. In total, there is one $60$ dimensional vector per frame. We scale all features to 0, 1 by max min scaling at participant level. Histograms of eye gaze and head pose are unimodal with different modes per participant, reflecting the differing setup of participant devices. We therefore zero mean the head pose and eye gaze vectors at participant level. We show a sample visualisation of OpenFace features in Figure 1. Head pose is depicted as a blue cube centred on the head pointing in the estimated direction of head pose. Eye gaze vectors are depicted in green. Facial landmarks are shown in red, though we only use a subset of 15 in this work. It is not straightforward to visualise FAUs so these are omitted.
@@
We establish the performance of the state-of-the-art voice activity projection (VAP) model Ekstedt and Skantze (2022b), an audio-only turn-taking model which has been used in several studies e.g. Ekstedt and Skantze (2022a); Ekstedt et al. (2023); Inoue et al. (2024). The VAP model is a transformer Vaswani (2017) based neural network trained to predict who will speak in the next 2 seconds: the training objective, shown in Figure 3. At each point in time there are four bins per speaker representing the next 2 seconds. The bins span the next 0-200, 200-600, 600-1200 and 1200-2000 milliseconds. If 50% of frames within a bin contain speech, we set the bin label to 1, otherwise the bin is labelled 0. This gives $2^{8}=256$ possible VAP states. The model has 5.8 M parameters, and full details are in Inoue et al. (2024); Ekstedt and Skantze (2022a).
@@
Like the VAP model, our model, multimodal VAP ($M$-VAP, Figure 4) consists of self- and cross-attention blocks. A self-attention block is NN stacked transformer decoder layers, where query ($q$), key ($k$) and value ($v$) are identical:
$\text{{Self-Attn}}(x)=\text{{Transformer}}\left(q=x,\ k=x,\ v=x\right)$
A transformer layer learns attention or the ’compatibility’ of output with the query via the key and value Vaswani (2017). In self-attention blocks, as $q=k=v$, the model learns temporal patterns in the input e.g. audio or video. In cross-attention blocks, we stack NN layers with two inputs $x_{1}$ and $x_{2}$. In each layer, we compute two transformer layers with shared weights:
@@
$\textsc{Bal. accuracy}=\frac{1}{2}\left(\frac{\#TP}{\#P}+\frac{\#TN}{\#N}\right)$ (3)
@@
We conduct an ablation study by training 4 different versions of the $M$-VAP model. We train and validate using the same procedure, but each of which receives audio and a subset of the visual features as input. We divide the visual features into gaze, head pose, facial action unit and facial landmarks groups (6, 6, 17, 30 dimensions receptively, excluding the scalar confidence score).
@@ 4.3 Multimodal turn-taking
We continue by investigating the performance of our proposed early and late fusion turn-taking models ($M$-VAP). We train on the Candor corpus (Section 3). For completeness, we also train a VAP model only using video features. This model is identical to the original VAP model with the removal of the audio feature extractor and a dimensionality of 60, reflecting the dimensionality of the visual feature vector.
@@
(FTO > +250 ms) v 0.72 $\downarrow$14% 0.82 $\downarrow$8% 0.47 $\downarrow$33% 68 $\downarrow$15%
a+v (e) 0.84 - 0.90 - 0.74 $\uparrow$5% 82 $\uparrow$3%
a+v (l) 0.86 $\uparrow$3% 0.90 $\uparrow$2% 0.74 $\uparrow$6% 83 $\uparrow$4%
@@
(FTO > +250 ms) v 0.70 $\downarrow$13% 0.80 $\downarrow$8% 0.45 $\downarrow$33% 66 $\downarrow$14%
a+v (e) 0.83 $\uparrow$2% 0.88 - 0.69 $\uparrow$4% 79 $\uparrow$3%
a+v (l) 0.83 $\uparrow$3% 0.89 $\uparrow$2% 0.70 $\uparrow$6% 80 $\uparrow$4%
@@
(FTO > 0 ms) v 0.78 $\downarrow$10% 0.87 $\downarrow$5% 0.43 $\downarrow$35% 70 $\downarrow$10%
a+v (e) 0.87 $\uparrow$2% 0.92 - 0.69 $\uparrow$6% 80 $\uparrow$3%
a+v (l) 0.87 $\uparrow$2% 0.92 - 0.71 $\uparrow$6% 83 $\uparrow$4%
@@
(FTO < -250 ms) v 0.72 $\downarrow$8% 0.82 $\downarrow$3% 0.42 $\downarrow$25% 64 $\downarrow$8%
a+v (e) 0.79 $\uparrow$3% 0.87 $\uparrow$2% 0.60 $\uparrow$7% 72 $\uparrow$3%
a+v (l) 0.80 $\uparrow$4% 0.87 $\uparrow$2% 0.62 $\uparrow$10% 74 $\uparrow$5%
@@
We report the average performance over 5 folds of the Candor corpus in Table 3. As before, we consider model performance 200 ms before a shift/hold (FTO > 250 ms). Our multimodal models significantly outperform the audio-only and video-only models ($p<0.01$). The late fusion model shows a 33% relative increase in the weighted $F_{1}$ score and a 66% relative increase in balanced accuracy over the audio-only model. The video-only model has the worst overall performance with a 33% relative reduction in the $F_{1}$ shift score.
@@
Next, we remove the requirement for a minimum period of silence between turns, increasing the number of shifts and holds considered (FTO > 0 vs FTO > 250 ms in Table 1). We also move the evaluation window to 200 ms before the end of a turn, ensuring the next speaker has not started to speak during evaluation. This hence captures the capacity of the model to predict upcoming shifts while the previous speaker is still speaking, akin to human turn-taking (Section 1). For this more comprehensive set of speaker transitions, our multimodal models outperforms both video-only and audio-only models. Best performance is achieved with the late fusion model, with a $6\%$ relative increase in the $F_{1}$ shift score and a $4\%$ relative increase in balanced accuracy (Table 3). We also evaluate hold/shift prediction before overlapping speech. Note that for the purposes of this evaluation, we exclude overlapping speech where there is no change in speaker 1 second after the overlap has concluded. This ensures that brief periods of overlap are excluded (i.e. backchannels). We find a 6% increase in balanced accuracy and a $10\%$ increase in the $F_{1}$ shift scores for the late-fusion $M$-VAP model (Table 3).
@@
The balanced accuracies in Tables 2 and 3 correspond with the 0 and 250 ms points in Figure 6. The gap in performance of Switchboard and Candor is consistent with our prior finding of a lower $F_{1}$ shift score on Switchboard (Table 2). Performance falls for both audio-only VAP and multimodal $M$-VAP models as the duration of silence between speaking turns increases, indicating that these turns are more challenging to predict. However, we find that the performance of the $M$-VAP models is significantly better than the audio-only VAP models across all FTOs ($p<0.01$ in all cases). Thus, the inclusion of visual information leads to improved performance across the complete range of holds and shifts in the corpus.
@@
Finally, we conduct an ablation study, training the late fusion version of the $M$-VAP model on audio and gaze, head pose, facial landmarks and facial action unit subsets. We report the 5-fold average performance in Table 4.
@@
(FTO > +250 ms) a + gaze 0.74 $\downarrow$14% 0.83 $\downarrow$8% 0.51 $\downarrow$31% 67 $\downarrow$20%
a + pose 0.84 $\downarrow$2% 0.89 - 0.72 $\downarrow$3% 80 $\downarrow$3%
a + lmks 0.79 $\downarrow$8% 0.86 $\downarrow$4% 0.62 $\downarrow$17% 74 $\downarrow$11%
a + faus 0.85 - 0.89 - 0.73 $\downarrow$2% 81 $\downarrow$2%
@@
(FTO > 250 ms) a + gaze 0.72 $\downarrow$13% 0.82 $\downarrow$8% 0.48 $\downarrow$35% 65 $\downarrow$19%
a + pose 0.82 - 0.87 - 0.68 $\downarrow$8% 78 $\downarrow$2%
a + lmks 0.78 $\downarrow$7% 0.85 $\downarrow$4% 0.59 $\downarrow$20% 72 $\downarrow$10%
a + faus 0.83 - 0.88 - 0.69 $\downarrow$6% 79 $\downarrow$2%
@@
(FTO > 0 ms) a + gaze 0.78 $\downarrow$10% 0.88 $\downarrow$5% 0.43 $\downarrow$39% 65 $\downarrow$20%
a + pose 0.87 - 0.92 - 0.67 $\downarrow$3% 79 $\downarrow$2%
a + lmks 0.83 $\downarrow$5% 0.90 $\downarrow$3% 0.57 $\downarrow$18% 73 $\downarrow$10%
a + faus 0.87 - 0.92 - 0.68 $\downarrow$2% 80 $\downarrow$4%
@@
(FTO < 250 ms) a + gaze 0.68 $\downarrow$15% 0.81 $\downarrow$7% 0.35 $\downarrow$44% 58 $\downarrow$22%
a + pose 0.79 $\downarrow$2% 0.86 - 0.58 $\downarrow$6% 71 $\downarrow$3%
a + lmks 0.74 $\downarrow$8% 0.84 $\downarrow$4% 0.49 $\downarrow$22% 66 $\downarrow$11%
a + faus 0.79 $\downarrow$2% 0.86 - 0.59 $\downarrow$4% 72 $\downarrow$2%
@@ 5 Discussion
Our re-implementation of the audio-only VAP PTTM Ekstedt and Skantze (2022b); Inoue et al. (2024) performed well on the Candor videoconferencing (VC) corpus. However, our new multimodal PTTM, $M$-VAP - which uses speech along with facial action units, landmarks, gaze, and head pose - significantly outperformed the VAP model. The performance increase was most notable on the $F_{1}$ shift metric, with a 6-10% relative increase in performance (Table 2). Returning to our first research question, our findings show that visual cues improve PTTM performance. This echoes the psycholinguistics literature underlining the importance of visual cues in turn-taking (Section 1).
@@
$M$-VAP has shown promising results, but it can be improved. OpenFace could be replaced by a more powerful learnable visual front-end, e.g. from the audio-visual speech recognition literature Fenghour et al. (2021). The model should be updated to handle audio and video at different sample rates (Section 3).